# Analogical Structure, Minimal Contextual Cues and Contrastive Distractors: Input Design for Sample-Efficient Linguistic Rule Induction

### v1

@@ 3.4 Error Analysis and Patterns
Figure 6: Error analysis across data organizations. Lightweight models (left) show robust performance with minimal errors ($3.3\%$) across all structures. LLMs (right, GPT-o3 zero-shot) show structure-dependent error patterns following our cognitive-inspired error taxonomy (Table 1). Shuffled structures particularly disrupt LLMs’ analogical reasoning, causing increased multi-constraint violations.
Figure 6 shows organizational structure affects error patterns across different model types. Lightweight models are robust ($errors<3.3\%$) across all organizational conditions, demonstrating consistent learning regardless of structural variations. LLMs show structure-dependent error patterns following our error taxonomy. Disrupted analogical structure (Shuffled condition) causes increased multi-constraint violations, with correct responses dropping from $55\%$ (Base) to $24\%$ (Shuffled) for gpt-o3. This shows that even reasoning-capable models rely heavily on structural organization for linguistic rule learning tasks.
@@ 3.5 Additional Analysis
Cross-Phenomenon Generalization
Figure 7: Cross-phenomenon validation using unspecified object alternations (bake-class verbs). Results show F1 performance using BERT+CNN architecture on Type-II data, averaged over 3 runs.
Figure 7 confirms Base >> Shuffled hierarchy replicates with bake-class verbs. This validates generalisability beyond causative constructions.
Figure 8: (A) Seen type vs. Unseen Type performance across training sizes.
Cross-Type Generalization
Figure 8 demonstrates successful rule abstraction: models achieve robust cross-type performance by 100200100-200 examples, though all conditions eventually converge to high performance with sufficient training data. This confirms learning of structural relationships rather than surface patterns.
@@ Discussion
Our systematic evaluation validates all three core claims. The finding that lightweight structured models ($F1=0.98$) exceed even gpt-o3’s reasoning capabilities ($F1=0.87$) demonstrates that systematic input organization can achieve superior linguistic rule learning with reduced computational requirements. The consistent organizational effects across both reasoning and standard LLMs suggest that analogical paradigm organization addresses fundamental challenges in pattern recognition that persist even in the most advanced current systems. The contrasting error patterns between model types suggest that organizational benefits operate through different mechanisms, with accelerated learning for lightweight models and improved error discrimination for LLMs. The cross-phenomenon and cross-type evaluation confirms that our approach captures general principles rather than construction-specific optimizations.
These findings indicate that structured input data organization is complementary to performance gains for linguistic rule learning beyond architectural advances or reasoning capabilities.
@@ 4 Related Work
Achieving competitive performance with reduced training needs remains a central challenge in NLP. Recent work emphasizes how reorganizing raw text into structured or hierarchically arranged contexts can significantly enhance model learning (liu2024enhancing; sourati2024arn). “Structured data” often refers to knowledge graphs, table data, databases (jiang-etal-2023-structgpt; li2023resdsql). liu2024enhancing instead reorganize text into hierarchical formats (e.g., Scope-Aspect-Description) to mimic human knowledge consolidation, thereby improving task performance in question-answering. Existing sample efficiency approaches in NLP include data selection (albalak2024a), data augmentation (kumar-etal-2020-syntax; dai-etal-2023-auggpt), curriculum learning (bengio2009curriculum; xu-etal-2020-curriculum), among others. However, they typically focus on scaling or sequencing existing data rather than fundamental reorganization principles. Our work instead targets linguistic rule learning through structure in a completion tasks inspired by cognitive principles.
@@
Analogical reasoning is the ability to perceive and use similarities on relational patterns rather than surface features (gentner2012analogical). Neural models have shown strong capabilities of analogical reasoning to generalize rules, such as word-level analogies (mikolov2013efficient; brown2020language; ushio-etal-2021-bert; webb2023emergent), concept analogies through analogical structure abduction (yuan-etal-2023-beneath), narrative-based or story-level analogies (nagarajah2022understanding; jiayang-etal-2023-storyanalogy; sourati2024arn), or analogical prompting where LLMs generate relevant exemplars or knowledge before solving the problem (yasunaga2024large). These approaches, like ours, suggest that explicit structural cues improve analogical processing.
@@
Contrastive learning has emerged as a powerful principle for neural representation learning. Systematic comparison between positive and negative examples enhances discriminative learning (chen2020simple; he2020momentum). However, existing contrastive approaches typically require large-scale training data and focus on representation learning rather than rule acquisition. Our systematic distractor design similarly implements a form of contrastive learning by organizing examples that emphasize critical distinctions between alternations. This enables models to learn from the structured contrasts between valid and invalid linguistic transformations.
@@
Neural approaches to linguistic rule learning are challenging despite recent advances. Neural language models demonstrate emergent syntactic capabilities (linzen-etal-2016-assessing; gulordava-etal-2018-colorless; hewitt-manning-2019-structural; mueller-etal-2024-context) but requires large-scale training. kann-etal-2019-verb examine alternations in transformer embeddings with limited success, thrush2020investigating find that systematic selectional preferences require careful training design. LLMs also struggle on metalinguistic tasks (thrush-etal-2024-strange). merlo-2023-blackbird introduces Blackbird Language Matrices (BLM) as structured evaluation paradigms for systematic rule assessment, enabling controlled testing of linguistic competence. However, they show limited efficiency in data-lean settings. Our work addresses these efficiency challenges through analogical paradigm organization, building on the BLM framework to achieve competitive argument structure learning with minimal training examples while demonstrating robustness across alternation types.
@@ 5 Conclusion
We investigated whether computational approaches inspired by cognitive principles can enable sample-efficient linguistic rule learning. Our analogical paradigm organization approach operationalises three organizational strategies, analogical structure, contrastive learning, and minimal contextual cues, through systematic input structuring rather than architectural scaling.
Our results demonstrate substantial sample efficiency gains across multiple evaluation dimensions. Lightweight models (BERT+CNN, $\sim 0.5M$ parameters) trained on only $100$ structured examples achieved $F1=0.95$, outperforming zero-shot GPT-o3 ($F1=0.87$) and stabilizing performance with just 100012001000-1200 training instances. Component analysis confirmed that each organizational principle contributes systematically, with analogical structure showing the largest effect. Cross-phenomenon validation with bake-class verbs and cross-type generalization experiments confirmed that benefits extend beyond construction-specific optimizations.
These findings have practical and theoretical implications. Practically, our approach enables competitive linguistic rule learning with much fewer resources than conventional LLM approaches. Theoretically, we demonstrate that systematic input organization can achieve efficiency gains that complement rather than require architectural scaling. This suggests that cognitive-inspired data structuring is a distinct optimization dimension.
Future work should explore cross-linguistic validation, integration with pre-training objectives, and automated structure discovery methods. This research shows that computational approaches inspired by cognitive principles can provide practical advances in sample-efficient linguistic rule learning.
@@ 6 Limitations
Our structured paradigms require expert linguistic knowledge to design appropriate analogical mappings and systematic distractors. Although we demonstrate cross-phenomenon generalisability with bake-class verbs, scaling this approach to diverse linguistic phenomena would require either linguistic expertise or automated methods for discovering optimal organizational structures.
@@
Our lightweight models rely on pre-trained encoders (BERT, RoBERTa, ELECTRA) that already encode linguistic knowledge from unstructured data. This creates attribution challenges in isolating whether performance gains stem purely from our organizational principles or from interactions with pre-existing linguistic representations. While our Base vs. Shuffled comparisons control for this confound, the fundamental attribution question remains.
@@
Our evaluation was exclusively on English verb alternations, a specific subset of linguistic phenomena at the syntax-semantics interface. Cross-linguistic validation and extension to other grammatical constructions (e.g., morphological patterns, syntactic transformations) are needed to establish broader applicability of our organizational principles.
@@
Our comparison between trained lightweight models and zero-shot LLMs may not reflect equivalent computational investments. While we demonstrate superior sample efficiency, a fairer comparison might involve fine-tuning LLMs on equivalent structured data or comparing against few-shot learning with similar computational budgets.
@@
Our approach targets pattern completion tasks that may favour analogical reasoning. Extension to diverse NLP applications (e.g., generation, reasoning) would provide stronger evidence for the general usability of our organizational principles.
@@
Our research presents minimal ethical concerns. Our synthetically generated data focuses on grammatical verb alternations rather than semantic content, minimizing bias propagation risks from LLM training data. Code, data, and experimental procedures will be made publicly available upon publication.
Appendix A Experimental Setup Details
A.1 Ablation Design Formalization
We systematically manipulate context structures to isolate organizational components. Given a full context matrix $C_{\text{full}}\in\mathbb{R}^{m\times n}$ and transformation matrix $T\in{0,1}^{m\times n}$, transformed contexts are computed as:
@@ A.1 Ablation Design Formalization
  • For Transposed Context, it simply involves taking the transpose of the full context matrix.
@@
A.2 Statistical Analysis Framework
@@
Structure (5 levels) $\times$ Model (8 levels) $\times$ Shots (3 levels) $\times$ CoT (2 levels) $\times$ Data Type (2 levels)
Generalization Gap Metric
$\text{GenGap}(t,s)=\frac{1}{|M|}\sum_{m\in M}\left[F1(m,s,\text{seen},t)-F1(m,s,\text{unseen},t)\right]$ (1)
where $t$ = training size, $s$ = structure type, $M$ = model set, and smaller gaps indicate better generalization.
Appendix B Linguistic Phenomena
B.1 Causative/Inchoative Alternation (Roll-Class Verbs)
The class of roll verbs, as categorised by Levin93, comprises verbs expressing dynamic actions with inherent motion characteristics. This class shows systematic alternation between:
  • Intransitive: Theme as subject (e.g., ‘The ball rolled”)
  • Transitive: Agent causes Theme motion (e.g.,‘ The player rolled the ball”)
Verb Inventory
B.2 Cross-Phenomenon Validation (Bake-Class Verbs)
  • Transitive: ‘The chef baked a cake”
  • Intransitive:‘ The chef baked” (object understood)
This tests whether organizational benefits generalize across alternation types rather than being phenomenon-specific optimizations.
@@ C.1 Encoder Comparison
Figure 9: F1 performance as a function of training size, comparing encoder architectures (BERT, RoBERTa, ELECTRA) with CNN architecture on Base structure. Train and test on type I data.
We employ BERT (bert-base-multilingual-cased) as our primary encoder to emphasize contributions attributable to data organization rather than encoder capabilities, as evidenced by Figure 9.
@@ C.2 Architecture Specifications
  • Function: Localizes sequential patterns for analogical mapping
@@
  • Function: Distributed pattern integration across entire context
@@ C.3 Training Configuration
Models are trained over $120$ epochs with learning rate $0.001$ using the Adam optimizer and batch size $100$. Early stopping with patience $10$ prevents overfitting. All models employ a max-margin objective using cosine similarity:
@@ D.1 Large Language Model Specifications
We evaluate eight advanced models across four families with varying parameter scales. This systematic comparison allows assessment of how organizational structure affects different model capacities and architectural approaches.
@@ D.2 Statistical Analysis Results
Complete multi-factor ANOVA results demonstrating significant main effects for Structure, Model, and Shots across all experimental conditions.
@@ D.3 Comprehensive Performance Results
Complete results across all models, structures, data types, and shot configurations with and without Chain-of-Thought reasoning.
@@
Appendix E Additional Analyses
E.1 Learning Dynamics Analysis
Figure 10 tracks error reduction as training progresses with the Base structure. ParadigmChange-RoleReplace errors show the most dramatic decrease, suggesting paradigm-level understanding develops earlier than fine-grained role distinctions. The Base structure facilitates more efficient error reduction across all types, providing better implicit feedback about linguistic rules consistent with error-driven learning theories.
E.2 Generalization Analysis
Figure 11: Generalization gap analysis for base and shuffled structures. Smaller gaps indicate better generalization. Base structure facilitates better transfer of linguistic rules to new contexts across all training sizes. Smoothing uses centered rolling window (window size=3) with confidence bands representing standard error propagation.
Figure 11 demonstrates that Base structure maintains consistent generalization patterns while Shuffled shows erratic performance in mid-range training sizes. At small training sizes, both structures show negative gaps, suggesting models avoid overfitting to surface lexical patterns. Base structure shows more stable and efficient generalization throughout the learning process.
E.3 Cross-Phenomenon Validation
Cross-phenomenon validation with bake-class verbs confirms that organizational benefits (Base >> Shuffled) replicate across different linguistic constructions, validating generalisability beyond causative alternations and demonstrating that our approach captures general organizational principles rather than construction-specific optimizations.

### v2

@@ 3.4 Error Analysis and Patterns
Figure 6: Error analysis across data organisations. Lightweight models (left) show robust performance with minimal errors ($3.3\%$) across all structures. LLMs (right, gpt-o3 zero-shot) show structure-dependent error patterns following our cognitive-inspired error taxonomy (Table 1). Shuffled structures particularly disrupt LLMs’ analogical reasoning, causing increased multi-constraint violations.
Figure 6 shows organisational structure affects error patterns across different model types. Lightweight models are robust ($errors<3.3\%$) across all organisational conditions, demonstrating consistent learning regardless of structural variations. LLMs show structure-dependent error patterns following our error taxonomy. Disrupted analogical structure (Shuffled condition) causes increased multi-constraint violations, with correct responses dropping from $55\%$ (Base) to $24\%$ (Shuffled) for gpt-o3. This shows that even reasoning-capable models rely heavily on structural organisation for linguistic rule learning tasks.
Figure 7 tracks error reduction as training progresses with the Base structure. ParadigmChange-RoleReplace errors show the most dramatic decrease, which suggests that paradigm-level information is learnt earlier than fine-grained role distinctions. The Base structure facilitates more efficient error reduction across all types, which provides better implicit feedback about linguistic rules, consistent with error-driven learning theories.
@@ 3.5 Additional Analysis
Cross-Phenomenon Generalisation
Figure 8: Cross-phenomenon validation using unspecified object alternations (bake-class verbs). Results show $F1$ performance using BERT+CNN architecture on Type II data, averaged over 3 runs.
Figure 8 shows that Base results are higher than Shuffled also with bake-class verbs. This suggests that our organisational benefits (Base >> Shuffled) generalise beyond causative alternations and indicates that our approach captures general organisational principles rather than construction-specific optimisations.
Figure 9: Seen type vs. Unseen Type performance across training sizes.
Cross-Type Generalisation
Figure 9 shows successful rule abstraction: models achieve robust cross-type performance by 100200100-200 examples, though all conditions eventually converge to high performance with sufficient training data. This confirms that structural relationships rather than surface patterns are being learnt.
Figure 10: Generalisation gap between Base and Shuffled structures. Smaller gaps indicate better generalisation. The Base structure facilitates better transfer of linguistic rules across training sizes.
Figure 10 shows that the Base maintains more stable and efficient generalisation throughout the learning process, while the Shuffled shows erratic performance in mid-range training sizes. At small training sizes, both structures show negative gaps, which suggests that the models avoid overfitting to surface lexical patterns. [^11] Details for the generalisation gap metric and smoothing method appear in Appendix B.
@@ Discussion
Our systematic evaluation validates all our three core claims. The finding that lightweight structured models ($F1=0.95$) exceed even gpt-o3 in zero-shot settings on our tasks ($F1=0.87$) demonstrates that systematic input organisation can achieve strong linguistic rule learning on controlled tasks with reduced task-specific data and computational requirements. The consistent organisational effects across both reasoning and standard LLMs indicate that analogical paradigm organisation addresses fundamental challenges in pattern recognition that persist even in the most advanced current systems. The contrasting error patterns between model types suggest that organisational benefits operate through different mechanisms, with accelerated learning for lightweight models and improved error discrimination for LLMs. The cross-phenomenon and cross-type evaluation confirms that our approach captures general principles rather than construction-specific optimisations.
These findings indicate that structured input data organisation is complementary to performance gains for linguistic rule learning beyond architectural advances or reasoning capabilities.
@@ 4 Related Work
Achieving competitive performance with reduced training needs remains a central challenge in NLP. Recent work emphasises how reorganising raw text into structured or hierarchically arranged contexts can significantly enhance model learning (Liu et al., 2024; Sourati et al., 2024). “Structured data” often refers to knowledge graphs, table data, databases (Jiang et al., 2023; Li et al., 2023). Liu et al. (2024) instead reorganise text into hierarchical formats (e.g., Scope-Aspect-Description) to mimic human knowledge consolidation, thereby improving task performance in question-answering. Existing sample efficiency approaches in NLP include data selection (Albalak et al., 2024), data augmentation (Kumar et al., 2020; Dai et al., 2023), curriculum learning (Bengio et al., 2009; Xu et al., 2020), among others. However, they typically focus on scaling or sequencing existing data rather than fundamental reorganisation principles. Our work instead targets linguistic rule learning through structure in a structured completion tasks inspired by cognitive principles.
@@
Analogical reasoning is the ability to perceive and use similarities on relational patterns rather than surface features (Gentner and Smith, 2012). Neural models have shown strong capabilities of analogical reasoning to generalise rules, such as word-level analogies (Mikolov et al., 2013; Brown et al., 2020; Ushio et al., 2021; Webb et al., 2023), concept analogies through analogical structure abduction (Yuan et al., 2023), narrative-based or story-level analogies (Nagarajah et al., 2022; Jiayang et al., 2023; Sourati et al., 2024), or analogical prompting where LLMs generate relevant exemplars or knowledge before solving the problem (Yasunaga et al., 2024). These approaches, like ours, suggest that explicit structural cues improve analogical processing.
In linguistics and cognitive science, work on argument structure and alternations has argued that children track families of related verbs and constructions when they learn causative alternations, rather than memorising isolated patterns. Bowerman (1974) studies how children acquire English causative verbs and the relation between cognitive, semantic, and syntactic development. Our approach is computationally inspired by this line of research, but it does not aim to model child learning directly. We instead study how analogical organisation of input, contrastive distractors, and minimal contextual cues affect sample efficiency and generalisation behaviour of neural models on controlled alternation tasks.
@@
Contrastive learning has emerged as a powerful principle for neural representation learning. Systematic comparison between positive and negative examples enhances discriminative learning (Chen et al., 2020; He et al., 2020). However, existing contrastive approaches typically require large-scale training data and focus on representation learning rather than rule acquisition. Our systematic distractor design similarly implements a form of contrastive learning by organising examples that emphasise critical distinctions between alternations. This enables models to learn from the structured contrasts between valid and invalid linguistic transformations.
@@
Neural approaches to linguistic rule learning are challenging despite recent advances. Neural language models demonstrate emergent syntactic capabilities (Linzen et al., 2016; Gulordava et al., 2018; Hewitt and Manning, 2019; Mueller et al., 2024) but require large-scale training. Kann et al. (2019) examine alternations in transformer embeddings with limited success. Thrush et al. (2020) find that systematic selectional preferences require careful training design. LLMs also struggle on metalinguistic tasks (Thrush et al., 2024). Merlo (2023) introduces Blackbird Language Matrices (BLM) as structured evaluation paradigms for systematic rule assessment, enabling controlled testing of linguistic competence. However, they show limited efficiency in data-lean settings. Our work addresses these efficiency challenges through analogical paradigm organisation, building on the BLM framework to achieve competitive argument structure learning with minimal training examples while demonstrating robustness across alternation types.
@@ 5 Conclusion
We investigated whether computational approaches inspired by cognitive principles can enable sample-efficient linguistic rule learning. Our analogical paradigm organisation approach operationalises three organisational strategies, analogical structure, contrastive learning, and minimal contextual cues, through systematic input structuring rather than architectural scaling.
Our results demonstrate substantial sample efficiency gains across multiple evaluation dimensions. Lightweight models (BERT+CNN, $\sim 0.5M$ parameters) trained on only $100$ structured examples achieved $F1=0.95$, and, in this setting, reached higher $F1$ than zero-shot GPT-o3 ($F1=0.87$) on the same structured tasks, and stabilising performance with just 100012001000-1200 training instances. Component analysis confirmed that each organisational principle contributes systematically, with analogical structure showing the largest effect. Cross-phenomenon validation with bake-class verbs and cross-type generalisation experiments confirmed that benefits extend beyond construction-specific optimisations.
These findings have practical and theoretical implications. Practically, our approach enables competitive linguistic rule learning with much fewer task-specific examples than conventional LLM-based setups in this controlled setting. Theoretically, we demonstrate that systematic input organisation can achieve efficiency gains that complement rather than require architectural scaling. This suggests that cognitive-inspired data structuring is a distinct optimisation dimension for linguistic rule induction.
Future work should explore cross-linguistic validation, integration with pretraining objectives, and automated structure discovery methods. Overall, our study shows that computational approaches inspired by cognitive principles can support sample-efficient linguistic rule learning and offer a complementary perspective on the behaviour of lightweight models and prompted LLMs.
@@ 6 Limitations
Our structured paradigms require expert linguistic knowledge to design appropriate analogical mappings and systematic distractors. In practice, we follow a hybrid process where experts select and validate seed sentences and design a small set of templates, and full paradigms are then generated automatically. Although we demonstrate cross-phenomenon generalisability with bake-class verbs, scaling this approach to diverse linguistic phenomena would require either linguistic expertise or automated methods for discovering optimal organisational structures.
@@
Our lightweight models rely on pre-trained encoders (BERT, RoBERTa, ELECTRA) that already encode linguistic knowledge from unstructured data. This creates attribution challenges in isolating whether performance gains stem purely from our organisational principles or from interactions with pre-existing linguistic representations. While our Base vs. Shuffled comparisons control for this confound, the fundamental attribution question remains.
@@
Our evaluation was exclusively on English verb alternations, a specific subset of linguistic phenomena at the syntax-semantics interface. Cross-linguistic validation and extension to other grammatical constructions (e.g., morphological patterns, syntactic transformations) are needed to establish broader applicability of our organisational principles.
@@
Our comparison between trained lightweight models and zero-shot and few-shot LLMs places the models in different learning regimes. The lightweight models are fine-tuned on 1001000100-1000 task-specific examples. The LLMs are evaluated in zero-shot and few-shot prompt-based settings without parameter updates. These settings are not matched in computational cost or data exposure. Our claims about lightweight models and LLMs are therefore restricted to this contrast and do not extend to general conclusions about LLM capabilities under full fine-tuning or matched training budgets.
@@
Our approach targets pattern completion tasks that may favour analogical reasoning. This controlled design allows us to isolate the effects of input organisation on linguistic rule induction, but it does not cover broad downstream applications. Extension to diverse NLP applications (e.g., generation, open-ended reasoning) would provide stronger evidence for the general usability of our organisational principles.
@@
Our research presents minimal ethical concerns. Our synthetically generated data focuses on grammatical verb alternations rather than semantic content, minimising bias propagation risks from LLM training data.
Acknowledgements
We thank the anonymous reviewers of ACL Rolling Review and EACL for their constructive feedback. We also thank colleagues, Vivi Nastase and Giuseppe Samo at Idiap Research Institute for valuable discussions. We gratefully acknowledge the partial support of this work by the Swiss National Science Foundation, through grant TMAG-1_209426 to PM.
Appendix A Linguistic Phenomena
A.1 Causative/Inchoative Alternation (roll-class verbs)
The class of roll verbs, as categorised by Levin (1993), comprises verbs expressing dynamic actions with inherent motion characteristics. This class shows systematic alternation between:
  • Intransitive: Theme as subject (e.g., “The ball rolled”)
  • Transitive: Agent causes Theme motion (e.g.,“The player rolled the ball”)
Verb Inventory
A.2 Cross-Phenomenon Validation (bake-class verbs)
  • Transitive: “The chef baked a cake
  • Intransitive:“The chef baked” (object understood)
Verb Inventory
The verbs within this class include: bake, carve, chop, clean, cook, crochet, draw, drink, dust, eat, embroider, hum, hunt, fish, iron, knead, knit, mend, milk, mow, nurse, pack, paint, play, plow, polish, read, recite, sew, sculpt, sing, sketch, sow, study, sweep, teach, type, vacuum, wash, weave, whittle and write.
This tests whether organisational benefits generalise across alternation types rather than being specific to a single alternation type.
Appendix B Experimental Setup Details
B.1 Ablation Design Formalisation
We systematically manipulate context structures to isolate organisational components. Given a full context matrix $C_{\text{full}}\in\mathbb{R}^{m\times n}$ and transformation matrix $T\in\{0,1\}^{m\times n}$, transformed contexts are computed as:
@@ B.1 Ablation Design Formalisation
  • For Transposed Context, it simply involves taking the transpose of the full context matrix:
@@
B.2 Statistical Analysis Framework
@@
We use a five-way ANOVA with factors: Structure (5 levels), Model (8 levels), Shots (3 levels), CoT (2 levels), and Data Type (2 levels).
Generalisation Gap Metric
We define the generalisation gap as
$\text{GenGap}(t,s)=\frac{1}{|M|}\sum_{m\in M}\bigl[F1_{m,s}^{\text{seen}}(t)-F1_{m,s}^{\text{unseen}}(t)\bigr]$ (1)
where $t$ is the training size, $s$ the structure type, $M$ the model set, and smaller gaps indicate better generalisation. Smoothing uses a centred rolling window ($size=3$) with confidence bands representing standard error propagation.
@@ C.1 Encoder Comparison
Figure 11: F1 performance as a function of training size, comparing encoder architectures (BERT, RoBERTa, ELECTRA) with the CNN architecture on Base structure. Train and test on Type I data.
We employ BERT (bert-base-multilingual-cased) as our primary encoder to emphasise contributions attributable to data organisation rather than encoder capabilities, as evidenced by Figure 11.
@@ C.2 Architecture Specifications
  • Function: localises sequential patterns for analogical mapping
@@
  • Function: distributed pattern integration across the entire context
@@ C.3 Training Configuration
Models are trained over $120$ epochs with learning rate $0.001$ using the Adam optimiser and batch size $100$. Early stopping with patience $10$ prevents overfitting. All models employ a max-margin objective using cosine similarity:
@@ D.1 Large Language Model Specifications
We evaluate eight advanced models across four families with varying parameter scales. This systematic comparison allows us to assess how organisational structure affects different model capacities and architectural approaches.
@@ D.2 Statistical Analysis Results
Complete multi-factor ANOVA results demonstrating significant main effects for structure, model, and number of shots across all experimental conditions.
@@ D.3 Comprehensive Performance Results
Complete results across all models, structures, data types, and number of shot configurations with and without Chain-of-Thought reasoning.